How much of other chemicals must be evaporated from 400 grams of a hand sanitizer that is 24% alcohol to strengthen it to a hand sanitizer that is 30% alcohol? Correct your answer to the nearest whole number.
Answer: grams
1
Expert's answer
2020-06-08T19:16:38-0400

400*0.24=96 - grams of alcohol

for 30%:

0.3(400-x)=96

400-x=320

x=80 grams